# Randomized methods for dynamical low-rank approximation

### v2

@@ Abstract
We introduce novel dynamical low-rank methods for solving large-scale matrix differential equations, motivated by algorithms from randomized numerical linear algebra. In terms of performance (cost and accuracy), our methods overperform existing dynamical low-rank techniques. Several applications to stiff differential equations demonstrate the robustness, accuracy and low variance of the new methods, despite their inherent randomness. Allowing augmentation of the range and corange, the new methods have a good potential for preserving critical physical quantities such as the energy, mass and momentum. Numerical experiments on the Vlasov-Poisson equation are particularly encouraging.
@@ 1 Introduction
for some vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ and time interval $[0,T]$. Examples of such differential equations include the discretization of partial differential equations in physics [EL18, CE22, CH22, Ein24], uncertainty quantification [BCSK17, AL24, FL18], quantum chemistry [SLC+24, CLS23] and also machine learning [SZK+22, SZCT23, SHNT23]; further applications will be provided in the next paragraphs. We refer to equation (1) as the full order model, and when the parameters $m$, $n$ are large, computing a dense solution to the problem becomes both computationally and memory intensive and is sometimes not feasible. To address this issue, reduced order modeling techniques have grown in popularity. A particularly successful one is the dynamical low-rank approximation [KL07, EL18], which uses low-rank compressions to efficiently store and evolve the solution over time.
On the other hand, the domain of randomized numerical linear algebra has emerged and is growing in popularity. A key achievement in the domain is the rangefinder algorithm, also called HMT, which comes with good theoretical guarantees [HMT11]. A direct application of the rangefinder is the randomized singular value decomposition (randomized SVD) algorithm which allows to approximate the truncated SVD of a (large) matrix $A\in\mathbb{R}^{m\times n}$ at a reduced cost. The randomized SVD algorithm scales well since it only requires a few matrix-vector multiplications, and such operations are particularly fast on GPUs. The generalized Nyström (GN) scheme [W+14, TYUC17, Nak20] is also a fast method for approximating the SVD of low-rank matrices, it is numerically stable and more accurate than the randomized SVD.
In this paper, we propose new time-stepping methods to approximate equation (1) with low-rank approximations. The new methods’ construction is inspired from the rangefinder, the randomized SVD and the generalized Nyström. At the moment, we are not aware of a direct formulation of randomized methods in the context time-dependent problems of the form (1), this paper is therefore an original attempt to link the domain of dynamical low-rank approximation and the domain of randomized numerical linear algebra. The new methods show a good potential for various applications in physics, uncertainty quantification, quantum chemistry, and machine learning. Before defining the new methods, we introduce notations and key results in the two domains.
@@ 1.1 Preliminary on randomized numerical linear algebra
The idea of using a randomized subspace iteration for finding the range goes back to [RST10]. Quickly after, the powered randomized rangefinder algorithm was proposed and analyzed in [HMT11], and the algorithm is now simply called rangefinder or HMT. This algorithm solves the problem of approximating the range of a matrix $A\in\mathbb{R}^{m\times n}$ with a prescribed target rank $r$ and oversampling parameter $p$. More precisely, it aims to find a matrix $Q\in\mathbb{R}^{m\times(r+p)}$ with orthogonal columns such that
@@
and numerical stability is achieved with an orthogonalization step after each application of $A$. The detailed procedure is given in [HMT11, Algorithm 4.4], and we briefly state here the main convergence theorem from [HMT11, Corrolary 10.10].
@@
In practice, the rangefinder is typically used with a relatively small oversampling parameter (typically $p\leq r$) and one or two power iterations are enough to get an average error close to the best error possible and also get a small variance. While random Gaussian matrices are generally a good choice, it appears that sub-sampled randomized trigonometric transforms (SRTTs) are slightly more accurate [MT20] and sparse maps allow faster matrix-vector multiplications.
@@
The randomized SVD scales particularly well since it only requires a few matrix-vector multiplications, and then costly operations are performed on small rectangular matrices. The truncation step is cost-free since the input is already given by its SVD, which can be truncated by keeping the $r$ first singular vectors/values. The analysis follows directly from Theorem 1.1, see also [HMT11, Theorem 10.5].
@@
Generalized Nyström (GN) schemes are an alternative to the randomized SVD algorithm. Let us introduce first Nyström’s scheme, originally proposed in [Nys30] and more recently popularized in [WS00]. Let $A\in\mathbb{R}^{n\times n}$ be a square positive semi-definite (PSD) matrix, that is $A\succeq 0$. Then, a rank-$r$ approximation is given by
@@ 1.2 Dynamical low-rank approximation
which is in fact a smooth manifold [LL03] embedded into $\mathbb{R}^{m\times n}$. The dynamical low-rank approximation is essentially the time-dependent variational principle (TDVP) applied to the vector field of the full order model and restricted to the above manifold of low-rank matrices; the DLRA evolves the projected dynamics
@@
$\left\|F(Y)-\mathcal{P}_{Y}\left[F(Y)\right]\right\|\leq\varepsilon_{r}.$
Note that the one-sided Lipschitz constant can be negative, e.g. on diffusive problems. Under these two assumptions, the solution $Y(t)$ to the DLRA compared to the solution $A(t)$ to the full order model verify the error bound
[…]
@@
The projector-splitting integrator [LO14] was the first integrator proposed, and its robust convergence was proved in [KLW16]. The projector-splitting integrator has a backward step and is therefore not suited for parabolic problems, then bases updates and Galerkin (BUG) methods were proposed in [CL22, CKL22, CKL24, CEKL24, Kus24]. Higher-order methods based on low-rank explicit Runge–Kutta schemes were proposed in [KV19, LCK24, NR25], and low-rank exponential Runge–Kutta methods for stiff differential equations are studied in [CV23]. Low-rank implicit schemes have also been studied in [NQE25, YEHS23, AC25, SV24, RV23], and a parallel-in-time low-rank integrator based on the parareal algorithm is discussed in [CGV23]. Dynamical CUR decomposition techniques have also been proposed in [DPN+23, NAB24, PN25], these methods have often less theoretical guarantees but have shown excellent results in practice and seem to be easier to implement and computationally less intensive than DLRA techniques, even though being closely related. Many of these methods have natural extensions to higher-dimensional problems with imposed tensor structures such as Tucker, tensor trains (or matrix product states) and tree tensor networks, see for example [KL10, LRSV13, LVW18, CL20, CL22, CLS23, GB24, EMP24, CKLS24].
The literature on applications of the dynamical low-rank approximation is now quite extensive, so we try to give a quick overview and more references can be found within the following papers. The DLRA has been successfully applied to various problems arising in quantum chemistry [SLC+24, CKLS24] and in physics: kinetics equations [EL18, Ein19, EOP20], radiation therapy [KS23], the Su–Olson problem [BEKK24], fission criticality [SEKM25] and also to stochastic differential equations [KNZ24, NT25, NT23]. Key properties of the numerical integrators in the context of physics are studied in [EJ21, EOS23, KEC23, EKS23]. Recent research also show applications in machine learning, such techniques are called dynamical low-rank training (DLRT) and are studied in [SZK+22, SZCT23, SL24, SZC+24, SYS25, KSW25].
@@ 1.3 Contributions and outline
As discussed in the previous section, the numerical issues due to the small singular values have been addressed by several numerical integrators. However, the low-rank integration of stiff differential equations is still challenging as they impose strong stepsize restrictions. Robust integrators based on exponential methods have been proposed in [OPW19, CV23], but they are limited to stiff differential equations with a Sylvester-like structure, and little is known for the low-rank integration of stiff differential equations with more general structures. The analysis of BUG-based methods [CL22, CKL24, Kus24] and projected Runge–Kutta methods [KV19] rely on the Lipschitz constant of the function of the vector field and have large hidden constants which are also observed numerically.
@@
The idea of bringing randomization to the dynamical low-rank approximation has been recently studied in [LCK24] where they propose randomized low-rank Runge–Kutta methods. The key idea of these methods is to use randomization (with the generalized Nyström formula) within explicit Runge–Kutta schemes to limit the rank growth in the stages. By doing so, they avoid the computation of projections onto the tangent space, which makes the schemes robust to the presence of small singular values, and in some cases more robust than the projected Runge–Kutta methods proposed in [KV19]. However, the approach is currently limited to explicit Runge–Kutta schemes and, as a consequence, the convergence depends again on the Lipschitz constant of the vector field, leading to large hidden constants when applied to stiff differential equations. In contrast to the projector-splitting and BUG integrators, low-rank explicit Runge–Kutta methods do not present a decomposition into substeps, which reduces the potential for parallelization and requires to implement these methods from scratch.
@@
Our main motivation is to develop BUG-like methods that can be applied to stiff differential equations without imposing strong stepsize restrictions. To achieve this, we introduce two new randomized time-stepping methods with a structure similar to existing BUG-based integrators. As a consequence, the new methods have a good potential for parallelization (in particular the dynamical generalized Nyström) and existing BUG implementations can be easily modified to implement the new methods. Additionally, we present novel rank-adaptive time-stepping methods where a Gaussian sampling is performed to estimate the errors at little extra cost. Our numerical experiments indicate that the new methods often outperform existing DLR integrators in terms of cost and accuracy, and demonstrate particular effectiveness in handling stiff differential equations characterized by large Lipschitz constants. With augmentation of the range and corange, the new methods also have a good potential for preserving critical physical quantities. A comprehensive convergence analysis of these methods remains an avenue for future research.
@@
The paper is organized as follows. In Section 2, we propose a new method that we call dynamical rangefinder, which can be used to estimate the range (or corange) of the solution of a matrix differential equation with a prescribed rank, and we also propose an alternative rank-adaptive method using a prescribed tolerance instead. We provide numerical experiments on a toy problem to show the accuracy of the estimated range, and we compare the new method with the rangefinder directly applied to the reference solution. Next, with the help of the new dynamical rangefinder, we propose in Section 3 two new time-stepping methods that we call respectively dynamical randomized SVD and dynamical generalized Nyström. We show that, when the solution to the full order model is exactly low-rank and that the estimated range contains the range of the full solution, then the new methods are exact. In Section 4, we apply the two new time-stepping methods to various problems on which we compare the performance (cost and accuracy) with different kind of DLRA techniques.
@@ 2 Dynamical randomized rangefinder
The notation ”$\min!$” is similar to the notation used in [FLLN24]. In this section, we start by discussing an idealized method where the exact solution to the full order model is available. Then, we propose a practical method to find such a matrix $Q_{h}$ without having to solve the full order model, and we propose a way to improve the approximation with dynamical power iterations. Finally, we discuss a method for automatically choosing the approximation rank for a prescribed tolerance using an error criterion with low probability of failure.
@@ 2.1 An idealized dynamical rangefinder method
    {.$&\accentset{\mbox{\large.}}{\mathcal{B}}(t)=F(A(t))\Omega,\quad t\in[0,h]\\ &\mathcal{B}(0)=A_{0}\Omega.\end{aligned}\right.$ (6)
@@
Fix a sketch matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$ and assume that $\operatorname{rank}(A(h))\leq r+p$. Then, the matrix $Q_{h}$ obtained from the above procedure verifies $\mathrm{Range}(Q_{h})=\mathrm{Range}(A(h))$ and therefore
@@
$\mathrm{Range}(Q_{h})=\mathrm{Range}(\mathcal{B}(h))=\mathrm{Range}(U_{h})=\mathrm{Range}(A(h)).$
As pointed out in the course of the above proof, the idealized procedure is, in fact, the rangefinder algorithm applied to the exact solution evaluated at time $h$, that is $A(h)$. It is also known that the projector-splitting integrator [LO14] and the BUG integrator [CL22] satisfy the exactness property on the exact flow.
@@ 2.2 A practical dynamical rangefinder with dynamical power iterations
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad t\in[0,h]\\ &B(0)=A_{0}\Omega.\end{aligned}\right.$ (7)
@@
    $&\accentset{\mbox{\large.}}{C}(t)=F(Q_{h}C(t))^{T}Q_{h},\quad t\in[0,h]\\ &C(0)=A_{0}^{T}Q_{h}.\end{aligned}\right.$ (8)
@@
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)W_{h}^{T})W_{h},\quad t\in[0,h]\\ &B(0)=A_{0}W_{h}.\end{aligned}\right.$ (9)

### v3

@@ Abstract
We introduce novel dynamical low-rank methods for solving large-scale matrix differential equations, motivated by algorithms from randomized numerical linear algebra. In terms of performance (ratio accuracy/cost), our methods can overperform existing dynamical low-rank techniques. Several applications to stiff differential equations demonstrate the robustness, accuracy and low variance of the new methods, despite their inherent randomness. Allowing augmentation of the range and corange, the new methods have a good potential for preserving critical physical quantities such as the energy, mass and momentum. Numerical experiments on the Vlasov-Poisson equation are particularly encouraging.
@@ 1 Introduction
for some vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ and time interval $[0,T]$. Examples of such differential equations include the discretization of partial differential equations in physics [EL18, CE22, CH22, Ein24], uncertainty quantification [BCSK17, AL24, FL18], quantum chemistry [SLC+24, CLS23] and also machine learning [SZK+22, SZCT23, SHNT23]; further applications will be provided in the next paragraphs. We refer to equation (1) as the full order model, and when the parameters $m$, $n$ are large, computing a dense solution to the problem becomes both computationally and memory intensive and is sometimes not feasible with current hardware. To address this issue, reduced order modeling techniques have grown in popularity. A particularly successful one is the dynamical low-rank approximation [KL07, EL18], which uses low-rank compressions to efficiently store and evolve the solution over time.
On the other hand, the domain of randomized numerical linear algebra has emerged and is growing in popularity. A key achievement in the domain is the rangefinder algorithm, also called HMT, which comes with good theoretical guarantees [HMT11]. A direct application of the rangefinder is the randomized singular value decomposition (randomized SVD) algorithm which allows to approximate the truncated SVD of a (large) matrix $A\in\mathbb{R}^{m\times n}$ at a reduced cost. The randomized SVD algorithm scales well since it only requires a few matrix-vector multiplications, and such operations are particularly fast on GPUs. The generalized Nyström (GN) scheme [W+14, TYUC17, Nak20] is also a fast method for approximating the SVD of low-rank matrices, it is numerically stable and also more accurate than the randomized SVD.
In this paper, we propose new time-stepping methods to approximate equation (1) by means of low-rank approximations. The new methods’ construction is inspired from the rangefinder, the randomized SVD and the generalized Nyström. At the moment, we are not aware of a direct formulation of randomized methods in the context time-dependent problems of the form (1), this paper is therefore an original attempt to link the domain of dynamical low-rank approximation and the domain of randomized numerical linear algebra. The new methods show a good potential for various applications in physics, uncertainty quantification, quantum chemistry, and machine learning. Before defining the new methods, we introduce notations and key results in the two domains.
@@ 1.1 Preliminary on randomized numerical linear algebra
The idea of using a randomized subspace iteration for finding the range goes back to [RST10]. Quickly after, the powered randomized rangefinder algorithm was proposed and analyzed in [HMT11], and the algorithm is now simply called rangefinder or HMT. This algorithm solves the problem of approximating the range of a matrix $A\in\mathbb{R}^{m\times n}$ with a prescribed target rank $r$ and oversampling parameter $p$. More precisely, the rangefinder computes a matrix $Q\in\mathbb{R}^{m\times(r+p)}$ with orthonormal columns such that
@@
and numerical stability is achieved with an QR step after each application of $A$. The detailed procedure is given in [HMT11, Algorithm 4.4], and we recall here the main convergence theorem from [HMT11, Corrolary 10.10].
@@
In practice, the rangefinder is typically used with a relatively small oversampling parameter, typically $p~\leq~r$. With power iterations, the constant in front of the best approximation error $\sigma_{r+1}$ contracts exponentially fast to one. While random Gaussian matrices are generally a good choice, it appears that sub-sampled randomized trigonometric transforms (SRTTs) are slightly more accurate [MT20]. Alternatively, sparse maps allow faster matrix-vector multiplications.
@@
The operator $\mathcal{T}_{r}:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ is the SVD truncated to rank $r$, also denoted $\mathcal{T}_{r}(A)=A_{[r]}$ in the literature. The randomized SVD scales particularly well since it only requires a few matrix-vector multiplications, and then costly operations are performed on small rectangular matrices. The truncation step is cost-free since the input is already given by its SVD, which can be truncated by keeping the $r$ first singular vectors/values. The analysis follows directly from Theorem 1.1, see also [HMT11, Theorem 10.5].
@@
Generalized Nyström (GN) schemes are an alternative to the randomized SVD algorithm. Let us introduce first Nyström’s scheme, originally proposed in [Nys30] and more recently popularized in [WS00]. Let $A~\in~\mathbb{R}^{n\times n}$ be a symmetric square positive semi-definite (SPSD) matrix. Then, a rank-$r$ approximation is given by
@@ 1.2 Dynamical low-rank approximation
which is in fact a smooth manifold [LL03] embedded into $\mathbb{R}^{m\times n}$. The dynamical low-rank approximation is essentially the time-dependent variational principle (TDVP) applied to the vector field of the full order model. Restricted to the above manifold of low-rank matrices, the DLRA evolves the projected dynamics
@@
$\left\|F(Y)-\mathcal{P}_{Y}\left[F(Y)\right]\right\|_{F}\leq\varepsilon_{r}.$
Note that the one-sided Lipschitz constant can be negative, e.g. on diffusive problems. Under these two assumptions, the solution $Y(t)$ to the DLRA compared to the full order model solution $A(t)$ verify the error bound
$\left\|A(t)-Y(t)\right\|_{F}\leq e^{t\ell}\left\|A_{0}-Y_{0}\right\|_{F}+\varepsilon_{r}\int_{0}^{t}e^{(t-s)\ell}ds,$ (5)
@@
The projector-splitting integrator [LO14] was the first integrator proposed, and its robust convergence was proved in [KLW16]. To overcome the limitations of the backward step in the projector-splitting integrator, bases updates and Galerkin (BUG) methods were proposed in [CL22, CKL22, CKL24, CEKL24, Kus24]. Higher-order methods based on low-rank explicit Runge–Kutta schemes were proposed in [KV19, LCK24, NR25], and low-rank exponential Runge–Kutta methods for stiff differential equations are studied in [CV23]. Low-rank implicit schemes have also been studied in [NQE25, YEHS23, AC25, SV24, RV23], and a parallel-in-time low-rank integrator based on the parareal algorithm is discussed in [CGV23]. Dynamical CUR decomposition techniques have also been proposed in [DPN+23, NAB24, PN25], these methods have often less theoretical guarantees but have shown excellent results in practice and seem to be easier to implement and computationally less intensive than DLRA techniques, even though being closely related. Many of these methods have natural extensions to higher-dimensional problems with imposed tensor structures such as Tucker, tensor trains (or matrix product states) and tree tensor networks, see for example [KL10, LRSV13, LVW18, CL20, CL22, CLS23, GB24, EMP24, CKLS24].
The literature on applications of the dynamical low-rank approximation is now quite extensive, so we try to give a quick overview and more references can be found within the following papers. The DLRA has been successfully applied to various problems arising in quantum chemistry [SLC+24, CKLS24] and in physics: kinetics equations [EL18, Ein19, EOP20], radiation therapy [KS23], the Su–Olson problem [BEKK24], fission criticality [SEKM25] and also to stochastic differential equations [KNZ24, NT25, NT23]. Key properties of the numerical integrators in the context of physics are studied in [EJ21, EOS23, KEC23, EKS23]. Recent research also show applications in machine learning, such techniques are called dynamical low-rank training (DLRT) and are studied in [SZK+22, SZCT23, SL24, SZC+24, SYS25, KSW25, BFN17, BFFN22].
@@ 1.3 Contributions and outline
As discussed in the previous section, the numerical issues due to the small singular values have been addressed by several numerical integrators. However, the low-rank integration of stiff differential equations is still challenging as they impose strong stepsize restrictions. Robust integrators based on exponential methods have been proposed in [OPW19, CV23], but they are limited to stiff differential equations with a Sylvester-like structure, and little is known for the low-rank integration of stiff differential equations with more general structures. The analysis of BUG-based methods [CL22, CKL24, Kus24] and projected Runge–Kutta methods [KV19] rely on the Lipschitz constant of the function of the vector field, implying large hidden constants which can be numerically observed.
@@
The idea of bringing randomization to the dynamical low-rank approximation has been recently studied in [LCK24] where they propose randomized low-rank Runge–Kutta methods. The key idea of these methods is to use randomization (with the generalized Nyström formula) within explicit Runge–Kutta schemes to limit the rank growth in the stages. By doing so, they avoid the computation of projections onto the tangent space, which makes the schemes robust to the presence of small singular values, and in some cases more robust than the projected Runge–Kutta methods proposed in [KV19]. However, the approach is currently limited to explicit Runge–Kutta schemes and, as a consequence, the convergence depends again on the Lipschitz constant of the vector field, leading to large hidden constants when applied to stiff differential equations. In contrast to the projector-splitting and BUG integrators, low-rank explicit Runge–Kutta methods do not present a decomposition into substeps, which reduces the potential for parallelization.
@@
Our main motivation is to develop BUG-like methods that can be applied to stiff differential equations without imposing strong stepsize restrictions. To achieve this, we introduce two new randomized time-stepping methods with a structure similar to existing BUG-based integrators. As a consequence, the new methods have a good potential for parallelization (in particular the dynamical generalized Nyström) and existing BUG implementations can be easily modified to implement the new methods. Additionally, we present novel rank-adaptive time-stepping methods where a Gaussian sampling is performed to estimate the errors at little extra cost. Our numerical experiments indicate that the new methods often outperform existing DLR integrators in terms of the ratio between accuracy and cost, and demonstrate particular effectiveness in handling stiff differential equations characterized by large Lipschitz constants. With augmentation of the range and corange, the new methods also have a good potential for preserving critical physical quantities. A comprehensive convergence analysis of these methods remains an avenue for future research.
@@
The paper is organized as follows. In Section 2, we propose a new method that we call dynamical rangefinder, which can be used to estimate the range (or corange) of the solution of a matrix differential equation with a prescribed rank, and we also propose an alternative rank-adaptive method using a prescribed tolerance instead. We provide numerical experiments on a toy problem to show the accuracy of the estimated range, and we compare the new method with the rangefinder directly applied to the reference solution. Next, we build on top of the new dynamical rangefinder, and we propose in Section 3 two new time-stepping methods that we respectively call dynamical randomized SVD and dynamical generalized Nyström. We show that, when the solution to the full order model is exactly low-rank and that the estimated range contains the range of the full solution, then the new methods are exact. In Section 4, we apply the two new time-stepping methods to various problems on which we compare the performance (ratio accuracy/cost) with different kind of DLRA techniques.
@@ 2 Dynamical randomized rangefinder
Here, the notation $\min!$” is similar to the notation used in [FLLN24]. In this section, we start by discussing an idealized method where the exact solution to the full order model is available. Then, we propose a practical method to find such a matrix $Q_{h}$ without having to solve the full order model, and we propose a way to improve the approximation with dynamical power iterations. Finally, we discuss a method for automatically choosing the approximation rank for a prescribed tolerance $\tau$ and probability of failure $\delta$.
@@ 2.1 An idealized dynamical rangefinder method
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{\mathcal{B}}(t)=F(A(t))\Omega,\quad t\in[0,h]\\ &\mathcal{B}(0)=A_{0}\Omega.\end{aligned}\right$. (6)
@@
Fix a sketch matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$ and assume that $\operatorname{rank}(A(h))\leq r+p$. Then, the matrix $\mathcal{Q}_{h}$ obtained from the above procedure verifies $\mathrm{Range}(\mathcal{Q}_{h})=\mathrm{Range}(A(h))$ and therefore
@@
$\mathrm{Range}(\mathcal{Q}_{h})=\mathrm{Range}(\mathcal{B}(h))=\mathrm{Range}(U_{h})=\mathrm{Range}(A(h)).$
As pointed out in the course of the above proof, the idealized procedure is, in fact, the rangefinder algorithm applied to the exact solution evaluated at time $h$, that is $A(h)$. The above property is similar to the exactness property of [LO14] and [CL22] on exact flows.
@@ 2.2 A practical dynamical rangefinder with dynamical power iterations
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad t\in[0,h]\\ &B(0)=A_{0}\Omega.\end{aligned}\right$. (7)
@@
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{C}(t)=F(Q_{h}C(t))^{T}Q_{h},\quad t\in[0,h]\\ &C(0)=A_{0}^{T}Q_{h}.\end{aligned}\right$. (8)
@@
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{B}(t)=F(B(t)W_{h}^{T})W_{h},\quad t\in[0,h]\\ &B(0)=A_{0}W_{h}.\end{aligned}\right$. (9)